OpenStudy (superhelp101):

When finding the sum of the arithmetic series 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means.

OpenStudy (vishweshshrimali5):

Please use a bracket for writing +(-2)

OpenStudy (superhelp101):

oh I'm sorry, the problem was written like that. I'll fix it! :) sryy

OpenStudy (vishweshshrimali5):

No need... Its good to see that you understand that :)

OpenStudy (superhelp101):

yes thank you for letting me know that! :)

OpenStudy (vishweshshrimali5):

Can you guess what is -3 here ? It is common difference of this AP

OpenStudy (superhelp101):

oh ok. So I Will be using the same formula?

OpenStudy (vishweshshrimali5):

Yep

OpenStudy (superhelp101):

okay, how do I find a_1?

OpenStudy (vishweshshrimali5):

a_1 is the first term of the AP. What is the first term here ?

OpenStudy (superhelp101):

4 is the first term

OpenStudy (vishweshshrimali5):

So it is a_1

OpenStudy (superhelp101):

so a_1=4....do I need to add all of them to get that total?

OpenStudy (vishweshshrimali5):

No use the formula for sum of an AP

OpenStudy (vishweshshrimali5):

\[\large{S_n = \cfrac{n}{2}[2a+(n-1)d]}\]

OpenStudy (vishweshshrimali5):

n = number of terms in the AP (Count them) d = You calculated before = -3 a = a_1 = first term = 4 S_n = sum the question wants

OpenStudy (superhelp101):

okay so n=6 d=need to find a=4 S_n= (do I add them)

OpenStudy (vishweshshrimali5):

n - correct d = a_2 - a_1 = second term - first term a = 4 S_n = Use the formula to calculate it

OpenStudy (superhelp101):

ohhhh okay, sry about that! I will calculate a let ya know in a sec... thank you in advance for your help!! :)

OpenStudy (vishweshshrimali5):

Your welcome (in advance) :D

OpenStudy (superhelp101):

;) :D

OpenStudy (superhelp101):

-117

OpenStudy (aum):

Although they want you to use the formula to calculate the arithmetic sum, with just six terms it is so much easier to simply add the 6 terms which can be done mentally while reading the problem!

OpenStudy (vishweshshrimali5):

How ?

OpenStudy (vishweshshrimali5):

@superhelp101 how ?

OpenStudy (superhelp101):

well I used the formula and just plugged in....@vishweshshrimali5 maybe I made an error...let me redo

OpenStudy (vishweshshrimali5):

d = -3 Thus: \[\large{S_n = 6/2(2*4 + (6-1)*(-3))}\] \[\large{\implies S_n = 3*(8 - 15)}\] \[\large{\implies S_n = -21}\]

OpenStudy (superhelp101):

oh yes....sry just looked back and that's what I got. What was aum taking about?

OpenStudy (vishweshshrimali5):

He was saying that since the number of terms (n) is small we can easily add up the terms. :) This method can be used to check the answer

OpenStudy (superhelp101):

oh ok

OpenStudy (vishweshshrimali5):

Also you can use this method if it is a multiple choice question. But, if it is a subjective question then don't use it :)

OpenStudy (superhelp101):

kk :)

OpenStudy (superhelp101):

so the sum is -21. And now I need to add them again?

OpenStudy (vishweshshrimali5):

No

OpenStudy (aum):

Just to double check your answer: The 4 + 1 cancels with -5 leaving you with -2 - 8 - 11 = -21

OpenStudy (vishweshshrimali5):

Add them and see if you get the same answer

OpenStudy (superhelp101):

ohhhhh I see makes sense! :)

OpenStudy (vishweshshrimali5):

What @aum is using a very intelligent method of solving the questions. But I would suggest you to use it ONLY when you have practised a large number of questions. After solving lot of questions, these type of things come in habit :)

OpenStudy (superhelp101):

okay, thank you for the tip! :) and thank you @aum for introducing it to me!

OpenStudy (vishweshshrimali5):

If you would notice, he very smartly observed that 4+1 = 5 and there is another -5 which would cancel out each other, leaving us only with 3 terms to add which is much easier than adding 6 terms.

OpenStudy (superhelp101):

yes, I see.. :)